Adrenaline is the hormone that triggers the release of extra glucose molecules in times of stress or emergency. A solution of 0.64 g of adrenaline in 36.0 g of CCl4 elevates the boiling point by 0.49 C. What is the molar mass of adrenaline if the Kb = 5.02 degrees Celsius/m?
1
Expert's answer
2013-02-08T03:53:43-0500
delta t = m x Kb x i
i = 1 so m = 0.49 / 5.02 = 0.09760 moles of adrenalin / kg of CCl4

also we have 0.64 g in 36.0 g of solvent or 0.64 x 1000 / 36 g of adrenalin /kg of solvent = 17.778 g

so in 1 liter of the solution we have 17.778 g = 0.09760 moles
1 mole of adrenalin has mass of 17.778 / 0.09760 = 182 g